OpenStudy (anonymous):

linear equation 6m-7=8m+19

OpenStudy (anonymous):

help me with steps please. ;)

OpenStudy (anonymous):

ohhh @misty1212

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@misty1212 @pooja195 @Michele_Laino

OpenStudy (anonymous):

Step 1: add 7 to both sides Step 2: subtract 8m from both sides Hint: you should now have: -2m = 26 Step 3: divide both sides by -2

OpenStudy (anonymous):

ok one second then while i figure it out

OpenStudy (phi):

to add 7 to both sides, write +7 on both sides like this 6m-7 + 7 =8m+19+7 notice you can simplify -7+7 on the left side. and 19+7 on the right side what do you get ?

OpenStudy (anonymous):

sorry about that 26

OpenStudy (anonymous):

but i thought you added 6m 8m first?

OpenStudy (phi):

you can do this in different order. but if we follow Xep's order (why not?) we do 6m-7 + 7 =8m+19+7 yes 19+7 is 26 so we have 6m -7 + 7 = 8m + 26 what about -7+7 ?

OpenStudy (anonymous):

sorry i'm kind of on and off. :P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

First you subtract 8m from both sides. Then add 7 to both sides. Then divide both sides of of the equation by -2 So ..... here is how to divide it. -2m / -2 = -26/-2 M = ?

OpenStudy (phi):

I assume you can simplify 7+ -7 , right ?

OpenStudy (anonymous):

0

OpenStudy (phi):

and 0 adding to anything (even 6m) does not change it. in other words 6m+0 is just 6m you now have 6m = 8m + 26 we want the "m"'s on the other side from the numbers one way is add -8m to both sides, like this -8m + 6m = -8m + 8m + 26

OpenStudy (phi):

on the right side you have 8 m's take away 8 m's (that is what -8m + 8m means) in other words , you get 0 for that part.

OpenStudy (phi):

on the left, it's 6 m's take away 8 m's (we will get a negative number, but that is ok)